# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ 3.3.1 Using cluster-period averages
$\bm{E}_{\mathrm{A}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{T}\bm{\Pi}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{A}}$ is the $J$-dimensional residual vector from the WLS fit of (9), $\bm{D}_{i,\mathrm{A}}$ is of dimensions $J\times J(J+1)$ with the $j$-th row being $\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ij}^{T}$, and $\bm{\Pi}_{i,\mathrm{A}}=\mathrm{diag}(\pi_{i1\cdot},\ldots,\pi_{iJ\cdot})$. Similar to $\overrightarrow{\mathbbm{1}}_{ijk}$, here, $\overrightarrow{\mathbbm{1}}_{ij}=(\mathbbm{1}_{ij,1}\ldots,\mathbbm{1}_{ij,J})^{T}$ with $\mathbbm{1}_{ij,j^{\prime}}=1$ if $j=j^{\prime}$ and $\mathbbm{1}_{ij,j^{\prime}}=0$ otherwise for $j^{\prime}=1,\ldots,J$. Then, the HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{A}}\bm{Q}(a,a^{\prime})^{T}$, and the HC variance estimator for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})\}$ is
@@
$\bm{E}_{\mathrm{A}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}}$ is the $J$-dimensional vector of residuals from the WLS fit of (10), $\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}$ is of dimensions $J\times J(1+p_{c})(J+1)$ with the $j$-th row being $(\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ij}^{T},\bm{G}_{i}(\mathcal{A})^{T}\otimes(\overrightarrow{\mathbbm{1}}_{ij}^{T}\otimes\bm{C}_{ij}^{c}))$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{A},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\times\allowbreak\bm{Q}(a,a^{\prime})^{T}$. Let $\bm{\gamma}_{j,\mathrm{A}}(a)$ denote the WLS coefficient vectors obtained using cluster-period averages if the full set of potential outcomes is available. That is, regress $\overline{Y}_{ij\cdot}(a)$ on $(1,\bm{C}_{ij}^{c})$ with weight $\pi_{ij\cdot}$ assuming $\overline{Y}_{ij\cdot}(a)$ is available for each cluster $i$ in period $j$. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ are given in Theorem 4.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{A}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,\allowbreak r_{iJ\cdot}(a))^{T}$. Under Assumptions 1 and 2, and regularity conditions (C1), (C4) and (C5) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}\allowbreak=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
Theorem 4 presents the consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$, and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, in the Löwner ordering. For the variance $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, the residual $r_{ij\cdot}(a)$ in Theorem 4 is in the form of $\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}$, which suggests that the optimal choice of $\bm{\gamma}$ is the coefficient of $I\pi_{ij\cdot}\bm{C}_{ij}^{c}$ in the OLS fit of $\widetilde{\epsilon}_{ij\cdot}(a)$ on $I\pi_{ij\cdot}\bm{C}_{ij}^{c}$, rather than $\bm{\gamma}_{j,\mathrm{A}}(a)$ from the WLS fit using cluster averages. Similar to the discussion following Theorem 2, $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ may not lead to better asymptotic efficiency compared to $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$. We omit the detailed discussion of $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$, the ANCOVA estimator obtained via regressing $\overline{Y}_{ij\cdot}$ on $(\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ij}^{T},\overrightarrow{\mathbbm{1}}_{ij}^{T}\otimes\bm{C}_{ij}^{c})$ with weights $\pi_{ij\cdot}$, because it is suboptimal and also may not improve the asymptotic efficiency of $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$. Similar to $\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})$ and $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$, $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$ is consistent and asymptotically normal for $\bm{\tau}(a,a^{\prime})$, and the associated HC variance estimator $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is conservative for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})\}$. We provide some clarifications on $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ regarding the special cases of $w_{ijk}=N_{ij}^{-1}$ and $w_{ijk}=1$ in Remark S2 in Section S2 of the Online Supplement.
@@ 3.3.2 Using scaled cluster-period totals
$\bm{E}_{\mathrm{T}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{T}\bm{\Pi}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{T}}$ is the $J$-dimensional vector of residuals from the OLS fit of (11), $\bm{D}_{i,\mathrm{T}}=\bm{D}_{i,\mathrm{A}}$, and $\bm{\Pi}_{i,\mathrm{T}}=\bm{I}_{J\times J}$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{T}}\bm{Q}(a,a^{\prime})^{T}$. The consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})$ and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}$ are given in Theorem 5.
@@
Define $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under Assumptions 1 and 2, and regularity conditions (C1) and (C6) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $I^{-2}\sum_{i=1}^{I}\widetilde{Y}_{ij\cdot}(a)^{4}=o(1)$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
$\bm{E}_{\mathrm{T}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}}$ is the $J$-dimensional vector of residuals from the OLS fit of (12), and $\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}=\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\times\allowbreak\bm{E}_{\mathrm{T},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}(a,a^{\prime})^{T}$. Let $\bm{\gamma}_{j,\mathrm{T}}(a)$ denote the OLS coefficient vectors obtained using scaled cluster-period totals if the full set of potential outcomes is available. That is, regress $\widetilde{Y}_{ij\cdot}(a)$ on $(1,\bm{C}_{ij}^{c})$ assuming $\widetilde{Y}_{ij\cdot}(a)$ is available for each cluster $i$ in period $j$. The consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})$ and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}$, in the Löwner ordering, are given in Theorem 6.

### v3

@@ 3.3.1 Using cluster-period averages
$\bm{E}_{\mathrm{A}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{\top}\bm{\Pi}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{A}}^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{A}}$ is the $J$-dimensional residual vector from the WLS fit of (9), $\bm{D}_{i,\mathrm{A}}$ is of dimensions $J\times J(J+1)$ with the $j$-th row being $\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ij}^{\top}$, and $\bm{\Pi}_{i,\mathrm{A}}=\mathrm{diag}(\pi_{i1\cdot},\ldots,\pi_{iJ\cdot})$. Similar to $\boldsymbol{\mathbbm{1}}_{ijk}$, here, $\boldsymbol{\mathbbm{1}}_{ij}=(\mathbbm{1}_{ij,1}\ldots,\mathbbm{1}_{ij,J})^{\top}$ with $\mathbbm{1}_{ij,j^{\prime}}=1$ if $j=j^{\prime}$ and $\mathbbm{1}_{ij,j^{\prime}}=0$ otherwise for $j^{\prime}=1,\ldots,J$. Then, the HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{A}}\bm{Q}(a,a^{\prime})^{\top}$, and the HC variance estimator for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})\}$ is
@@
$\bm{E}_{\mathrm{A}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{A}}\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{A}}\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{A}}^{\mathrm{adj}}$ is the $J$-dimensional vector of residuals from the WLS fit of (10), $\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}$ is of dimensions $J\times J(1+p_{c})(J+1)$ with the $j$-th row being $(\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ij}^{\top},\bm{G}_{i}(\mathcal{A})^{\top}\otimes(\boldsymbol{\mathbbm{1}}_{ij}^{\top}\otimes\bm{C}_{ij}^{c}))$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{A},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\times\allowbreak\bm{Q}(a,a^{\prime})^{\top}$. Let $\bm{\gamma}_{j,\mathrm{A}}(a)$ denote the WLS coefficient vectors obtained using cluster-period averages if the full set of potential outcomes is available. That is, regress $\overline{Y}_{ij\cdot}(a)$ on $(1,\bm{C}_{ij}^{c})$ with weight $\pi_{ij\cdot}$ assuming $\overline{Y}_{ij\cdot}(a)$ is available for each cluster $i$ in period $j$. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$ are given in Theorem 4.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}_{j,\mathrm{A}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,\allowbreak r_{iJ\cdot}(a))^{\top}$. Under Assumptions 1 and 2, and regularity conditions (C1), (C4), and (C5) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}\allowbreak=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
Theorem 4 presents the consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$, and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, in the Löwner ordering. For the variance $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})\}$, the residual $r_{ij\cdot}(a)$ in Theorem 4 is in the form of $\widetilde{\epsilon}_{ij\cdot}(a)-I\pi_{ij\cdot}\bm{C}_{ij}^{c}\bm{\gamma}$, which suggests that the optimal choice of $\bm{\gamma}$ is the coefficient of $I\pi_{ij\cdot}\bm{C}_{ij}^{c}$ in the OLS fit of $\widetilde{\epsilon}_{ij\cdot}(a)$ on $I\pi_{ij\cdot}\bm{C}_{ij}^{c}$, rather than $\bm{\gamma}_{j,\mathrm{A}}(a)$ from the WLS fit using cluster averages. Similar to the discussion following Theorem 2, $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ may not lead to better asymptotic efficiency compared to $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$. We omit the detailed discussion of $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$, the ANCOVA estimator obtained via regressing $\overline{Y}_{ij\cdot}$ on $(\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ij}^{\top},\boldsymbol{\mathbbm{1}}_{ij}^{\top}\otimes\bm{C}_{ij}^{c})$ with weights $\pi_{ij\cdot}$, because it is suboptimal and also may not improve the asymptotic efficiency of $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$. Similar to $\widehat{\bm{\tau}}_{\mathrm{A}}(a,a^{\prime})$ and $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$, $\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})$ is consistent and asymptotically normal for $\bm{\tau}(a,a^{\prime})$, and the associated HC variance estimator $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is conservative for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{A}}^{\mathrm{ancova}}(a,a^{\prime})\}$. We provide some clarifications on $\widehat{\tau}_{j,\mathrm{A}}(a,a^{\prime})$ and $\widehat{\tau}_{j,\mathrm{A}}^{\mathrm{adj}}(a,a^{\prime})$ regarding the special cases of $w_{ijk}=N_{ij}^{-1}$ and $w_{ijk}=1$ in Remark S3 in Section S2 of the Online Supplement.
@@ 3.3.2 Using scaled cluster-period totals
$\bm{E}_{\mathrm{T}}=\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)^{-1}\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{\top}\bm{\Pi}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)\left(\sum_{i=1}^{I}\bm{D}_{i,\mathrm{T}}^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{T}}$ is the $J$-dimensional vector of residuals from the OLS fit of (11), $\bm{D}_{i,\mathrm{T}}=\bm{D}_{i,\mathrm{A}}$, and $\bm{\Pi}_{i,\mathrm{T}}=\bm{I}_{J\times J}$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{T}}\bm{Q}(a,a^{\prime})^{\top}$. The consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})$ and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}$ are given in Theorem 5.
@@
Define $r_{ij\cdot}(a)=\widetilde{Y}_{ij\cdot}(a)-\overline{Y}_{\cdot j\cdot}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{\top}$. Under Assumptions 1 and 2, and regularity conditions (C1) and (C6) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $I^{-2}\sum_{i=1}^{I}\widetilde{Y}_{ij\cdot}(a)^{4}=o(1)$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
$\bm{E}_{\mathrm{T}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{T}}\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{T}}\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{T}}^{\mathrm{adj}}$ is the $J$-dimensional vector of residuals from the OLS fit of (12), and $\bm{D}_{i,\mathrm{T}}^{\mathrm{adj}}=\bm{D}_{i,\mathrm{A}}^{\mathrm{adj}}$. The HC variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\times\allowbreak\bm{E}_{\mathrm{T},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}(a,a^{\prime})^{\top}$. Let $\bm{\gamma}_{j,\mathrm{T}}(a)$ denote the OLS coefficient vectors obtained using scaled cluster-period totals if the full set of potential outcomes is available. That is, regress $\widetilde{Y}_{ij\cdot}(a)$ on $(1,\bm{C}_{ij}^{c})$ assuming $\widetilde{Y}_{ij\cdot}(a)$ is available for each cluster $i$ in period $j$. The consistency and asymptotic normality of $\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})$ and the conservativeness of $\widehat{\mathbf{Cov}}_{\mathrm{HC}}\{\widehat{\bm{\tau}}_{\mathrm{T}}^{\mathrm{adj}}(a,a^{\prime})\}$, in the Löwner ordering, are given in Theorem 6.